OpenStudy (anonymous):

Three vertices of square WXYZ are W(1, 4), X(4, 0), and Y(0, –3). What are the coordinates of the fourth vertex?

OpenStudy (anonymous):

Will star and medal!

OpenStudy (anonymous):

mid point of diagonals is same.

OpenStudy (anonymous):



OpenStudy (landshark5):

Try putting it on a graph. It will help you visualize it better.

OpenStudy (anonymous):

Thank you guys!

OpenStudy (anonymous):

\[\frac{ 4+x }{ 2 }=\frac{ 1+0 }{ 2 },find~ x\] \[\frac{ y+0 }{ 2 }=\frac{ 4+(-3) }{ 2 }\] find y